OpenStudy (anonymous):

Find values of a and b so that y = a·b^x and the line y = x + 4 are tangent at x = 0. I have no idea what to do here. It wasn't on the written homework due in class today and the teacher didn't go over it, an apparently this is going to be on the test tomorrow, so I need this broken down pretty well so I can memorize the steps for tomorrow. Please and thank you!

OpenStudy (anonymous):

is this caclulus?

OpenStudy (anonymous):

*calculus?

OpenStudy (anonymous):

not sure how to proceed without it on thing you know is that they have to meet at \(x=0\) so \[ab^0=0+4\] in other words \[a=4\]

OpenStudy (anonymous):

Yes, it is calculus.

OpenStudy (anonymous):

ok well we got \(a\) right?

OpenStudy (anonymous):

clear or no? then we can find \(b\)

OpenStudy (anonymous):

Yes, I understand how to get a now, I think. I'll do more practice problems later to make sure. What about b?

OpenStudy (anonymous):

i am thinking the answer i am getting is weird hold on a sec

OpenStudy (anonymous):

ok i think we can to it

OpenStudy (anonymous):

do you know what the derivative of \(b^x\) is in general?

OpenStudy (anonymous):

The derivative of b^x is (b^x)*ln(b).

OpenStudy (anonymous):

ok so the derivative of \[4b^x\] is \[4\ln(b)b^x\]

OpenStudy (anonymous):

the slope of \(y=x+4\) is \(1\) that means you want the derivative to be 1 at \(x=0\)

OpenStudy (anonymous):

set \[4\ln(b)=1\] ( that is the derivative at 0) and solve for \(b\)

OpenStudy (anonymous):

clear how to do that, or no?

OpenStudy (anonymous):

I think so. The answer I got was \[\sqrt[4]{e}\] Does that sound about right? Also, sorry for the equation being on a different line, I don't know how to get it to stay in the same line as the first sentence.

OpenStudy (anonymous):

no problem yeah or \[b=e^{\frac{1}{4}}\]same thing

OpenStudy (anonymous):

making your function \[f(x)=4e^{\frac{x}{4}}\]

OpenStudy (anonymous):

looks right, check here http://www.wolframalpha.com/input/?i=plot+4e^%28x%2F4%29,+x%2B4

OpenStudy (anonymous):

Oh okay. Yeah, that was the correct value for b according to Webstudy. I really appreciate the help, thank you! That problem really confused me.

OpenStudy (anonymous):

no problem, glad to help